# Stable Score Distillation for High-Quality 3D Generation

### 14 Dec 2023 v1

@@ 5.3 Evaluation on Text-to-3D Generation
Comparison with SDS.
Figure 6: SSD produces finer details and more plausible colors compared to SDS. See more comparisons in App. Fig. 18.
SDS and SSD are compared on text-to-3D generation, as shown in Fig. 6. The corresponding prompts are provided in App. Sec. 8. We show that SSD not only generates finer local features, exemplified by the coat color of the rabbit and plate textures in (a), and the eye region details of the robot in (c), but also alleviates the color-saturation problem as demonstrated by the mouth region of the cat in (b), and the body color of the robot in (c).
High-quality 3D Content Generation. We evaluate the capability of SSD with diverse text prompts. As shown in Fig. 1 and Fig. 2, SSD is compatible with different 3D representations and is able to generate high-quality 3D Gaussian Splatting/NeRF objects and diverse effects. More qualitative results can be found in App. Sec. 9 with more objects, effects and avatars.
User Studies. We compare SSD to SDS with user studies on preferences. Participants are presented with two videos side by side, both generated from the same text prompt. We ask the users to evaluate on two aspects: a) color authenticity (existence of over-saturation) and b) details (existence of over-smoothness). We randomly select 100 prompts from the DreamFusion gallery for evaluation, with each prompt assessed by 10 different users, resulting in a total of 1,000 pairwise comparisons. The result reveals a preference of 3D models generated with our method over SDS, $83.2\%$ for more realistic color and $71.4\%$ for finer details.
@@ 5.4 Ablation Study
The proposed Stable Score Distillation (SSD) introduces a hyper-parameter $M$ which denotes the timestep smaller than which we utilize the mode-seeking term instead of the mode-disengaging term for generation guidance. Note that setting $M=0$ results in the same formulation as concurrent work of CSD [49]. Fig. 7 shows that a recommended rule of thumb for setting $M$ is $200$. Reducing SSD to utilize the mode-disengaging term alone with $M=0$ results in 3D content with severe color saturation and excessive floaters. The corresponding prompts are provided in App. Sec. 8.
Figure 7: Ablation on timestep threshold $M$ with 3D generation.
6 Discussion and Conclusion
Limitation. Although our proposed method can generate high-quality 3D assets, due to the lack of proper constraints for multi-view consistency as stated in Sec. 4.1, resolving the Janus problem is beyond the scope of this study.
Societal Impact. Utilizing Stable Diffusion (SD) as the 2D prior for generating 3D content raises the possibility of inheriting societal biases embedded in the loosely curated web data used for SD training. We emphasize the importance of ethical and non-offensive usage for this technique.
Conclusion. We propose Stable Score Distillation (SSD) for high-quality 3D content generation, based on a comprehensive understanding of the widely used Score Distillation Sampling (SDS). We interpret SDS as a combination of mode-disengaging, mode-seeking and variance-reducing terms, analyzing their distinct properties. This interpretation allows us to harness each term to its fullest potential and to leverage their complementary nature. Our analysis establishes rich connections to prevalent observations and practices in 3D content generation. Extensive experiments demonstrate the effectiveness of our approach for generating high-fidelity 3D content without succumbing to issues such as over-smoothness and over-saturation. Notably, our method performs well even with small Classifier-Free Guidance (CFG) scale and NeRF representation alone.
@@ 6 Discussion and Conclusion
In this supplementary material, we provide more implementation details in Sec. 7. Sec. 8 provides the prompts used for generating the results in the main paper. Sec. 9 presents more qualitative results of SSD on diverse prompts. Sec. 10 gives more ablation and analysis on the design choices of our method. We provide proofs in Sec. 11 and Sec. 12, as well as numerical studies in Sec. 13 to support our method’s assumptions and analysis.
7 More Implementation Details
On threestudio, our experiments are based on the ProlificDreamer [45] implementation from the framework [9], with several slight modifications besides those discussed in Sec. 5.1. First, we utilize the progressive hash grid implemented in threestudio as we find it beneficial to the smoothness of geometry and the generation success rate. Additionally we apply the orientation loss whose strength grows linearly from 10.0 to 1000.0 in 10000 steps, together with sparsity loss whose strength is 10.0. We do not apply opaque loss throughout the experiments. We train each high-resolution 3D asset with 15000 steps, with the first 10000 trained with $64\times 64$ resolution and the other 5000 trained with resolution $512\times 512$.
On LucidDreamer’s codebase, we mainly adopt their provided configuration files.
8 Prompts Gallery
The prompts used for generating the 2D results in Fig. 4 are “a bunny on a pancake” and “a hamburger”.
The prompts used for 3D generation in Fig. 6 are “a zoomed out DSLR photo of a baby bunny sitting on top of a pile of pancake”, “a cat with a mullet”, and “a DSLR photo of a humanoid robot using a laptop”, respectively.
The prompts for generating the 3D results in Fig. 7, presented from left to right, are “a cat with a mullet”, “a DSLR photo of A DMC Delorean car”, and “a 20-sided die made out of glass”, respectively.
@@ 9 More Qualitative Results
In this section, we provide more text-to-3D (NeRF) generation results with diverse objects, effects and avatars.
High-quality text-to-3D generation. We present additional 3D generation results with more prompts in Fig. 11, Fig. 12 and Fig. 13.
Comparison with SDS on text-to-avatar generation. To assess the efficacy of the proposed SSD for avatar generation, we employ the text-to-avatar generation framework DreamWaltz [14] and contrast the standard SDS optimization supervision with our proposed SSD. As illustrated in Fig. 8, the utilization of SSD results in enhanced generation details (a-b) and heightened plausibility (c).
@@
Figure 10: Ablation on timestep threshold $M$ with 2D generation. The prompt is “a bunny on top of a pancake”. As $M$ increases, the generated results become smoother, exhibiting less implausible colors and textures, in sacrifice of the delicacy of local features. We use $M=200$ in our experiments to strike a balance between the smoothness and fineness of local features.
@@
Figure 13: More qualitative results on text-to-NeRF generation.
10 More Ablation
Verifying the efficacy of our adaptive variance reduction. Here we validate the design of our adaptive variance-reduction scheme proposed in Sec. 4.4.2 by showing that it is a general technique. Since SSD only activates the variance-reduction scheme when $t<200$, which amounts to affecting only 1/5 of the total training signals, we conduct a stronger ablation study to make the demonstration more credible. Specifically, we transplant our variance-reduction scheme to DreamFusion, and compare the quality of the generated 3D assets with and without the proposed variance reduction formulation. The comparisons are presented in Fig. 14, Fig. 15 and Fig. 16. The results show that the adaptive variance-reduction scheme is able to generalize to benefit SDS. Note that the random seed is fixed throughout.
Comparison with SDS and CSD [49] on text-to-3D generation. Note that CSD is a special case of SSD when $M=0$. In Fig. 18, we compare the quality of NeRF generations of CSD, SDS and SSD. In summary, CSD suffers from excessive floaters, and the die and the crab generated by CSD exhibit implausible colors. SDS, on the other hand, demonstrates higher color plausibility but is prone to over-smoothness. In contrast, our SSD combines the merits from both worlds, yielding intricate 3D assets with realistic coloration. The shortcomings of CSD and SDS are attributed to the factors elucidated in Sec. 4.6.
Ablation on $M$. We ablate on $M$ with a 2D experiment. We show in Fig. 10 that $M=200$ decently balances the color plausibility, smoothness and fineness of local features. We use $M=200$ in SSD throughout our experiments.
Ablation on utilization of the mode-seeking term. We identify the mode-disengaging term as the primary training signal. As illustrated in Fig. 9, the 3D asset generated with the mode-seeking term alone is extremely smooth without any delicate local features. In contrast, utilizing the mode-disengaging term alone can result in reasonable 3D generations as shown in Fig. 18 (left).
Validating the correlation between implausible local features and $t$. We conduct a 2D experiment in which only the mode-disengaging term takes effect, and we modify the upper limit and lower limit of $t$, recording the 2D generations. In Fig. 17 it is evident that training signals from small $t$ incur implausible local features. We attribute the problem to reasons discussed in Sec. 4.2.
Figure 14: Comparison on prompt “a 20-sided die made out of glass”. Note that our proposed adaptive variance reduction scheme accelerates the learning process of DreamFusion. Moreover, the resulting 3D asset exhibits more refined local details, as evidenced by the clarity of digits. Additionally, the geometry appears more plausible, with clearer edges between facets, underscoring the effectiveness of the adaptive variance reduction technique.
Figure 15: Comparison on prompt “a baby dragon drinking boba”. With the proposed adaptive variance reduction scheme, DreamFusion generates clear outline of the boba with only 2000 training steps. The accelerated learning pace on local features facilitates the successful generation of detailed boba, in contrast to the fuzziness observed in the absence of the adaptive variance reduction scheme.
Figure 16: Comparison on prompt “a bald eagle carved out of wood”. With the adaptive variance reduction scheme, DreamFusion learns faster and generates plausible squama and colors of the carved eagle with only 3000 training steps. This scheme effectively alleviates the over-smoothing problem and the final generation shows delicate squama whereas the same region is overly smooth without the scheme.
@@ 10 More Ablation
11 Proof for Closed Form Solution of c
@@ 13 More Numerical Experiments
Figure 18: Comparison between CSD, SDS and SSD on text-to-NeRF generation. From top to bottom, the prompts are “a 20-sided die made out of glass”, “a bunch of colorful marbles spilling out of a red velvet bag”, “a crab, low poly”, “a DSLR photo of a cake covered in colorful frosting with a slice being taken out, high resolution”, and “a DSLR photo of edible typewriter made out of vegetables”.
@@
Figure 19: $||\hat{\boldsymbol{\epsilon}}_{\phi}||$ increases as diffusion timestep $t$ increases. This together with Fig. 21 justifies our design of adaptive variance reduction.
Figure 20: The optimal $c(\boldsymbol{x},t)$ lies within the extremes of our proposed proxy $r$. For each timestep, we compute $r$ as defined in Eq. 8 and $c$ as defined in Eq. 7. Note that with the 8192 samples as discussed in Sec. 13, $c$ is unique while $r$ can take 8192 different values. Therefore the maximum and minimum of these $r$ are visualized separately. It is evident that $r$ and $c$ are numerically close with similar trend, validating that $r$ is a feasible proxy for $c$.
@@
Figure 21: Linear correlation [31] between $\hat{\boldsymbol{\epsilon}}_{\phi}$ and $\boldsymbol{\epsilon}$ increases with diffusion timestep $t$. This together with Fig. 19 justifies our design of adaptive variance reduction.
@@
Figure 22: Comparison between norms of the mode-seeking and mode-disengaging terms. Note that the y-axis is plotted in log scale. $||\hat{\boldsymbol{\epsilon}}_{\phi}||\in[80,130]$ while $||h||\in[2,5.5]$. The vast norm difference between the two terms constitutes an important justification for our estimator-rescaling in the proposed stable score distillation (SSD) Algorithm 1 (L15).

### 07 Feb 2024 v2

@@ 5.3 Evaluation on text-to-3D Generation
Comparison with baselines. We compare SSD with SOTA methods: DreamFusion [24], Fantasia3D [3], Magic3D [17] and ProlificDreamer [39] on text-to-3D generation and provide the results in Fig. 4 and Fig. 5. SSD generates results that are more aligned with the prompts (e.g., the plate generation in the cookie example), plausible in geometry (e.g., the plush dragon toy example) and color (e.g., compared to Magic3D), and delicate.
High-quality 3D Content Generation. We evaluate the capability of SSD with diverse text prompts. As shown in Fig. 2 and Fig. 1, SSD is able to generate high-quality general 3D objects and diverse effects. With different 3D representations, it is highly efficient.
Compatibility with frameworks. In Fig. 3 and Fig. 8, we show SSD’s compatibility with existing 3D generation frameworks, no matter if they are general-purpose or specialized. SSD can be readily incorporated into them for quality improvement.
User Studies. We compare SSD to the four outlined baselines with user preference studies. Participants are presented with five videos side by side, all generated from the same text prompt. We ask the users to evaluate on two aspects: a) plausibility and b) details (existence of over-smoothness). We randomly select 100 prompts from the DreamFusion gallery for evaluation, with each prompt assessed by 10 different users, resulting in a total of 1,000 independent comparisons. The result reveals a preference of 3D models generated with our method over other baselines, $53.2\%$ for more generation plausibility and $61.4\%$ for finer details.
@@ 5.4 Ablation Study
In Fig. 6, Fig. 13, and Fig. 14, we assess all the designs of SSD. Removing of any of them can result to a general decrease in generation quality. Specifically, we see higher $M$ causes the result to be more smooth with fewer local details, as expected. The three terms, namely mode-seeking, mode-disengaging and variance reduction terms are all necessary for successful 3D generation. The variance reduction term can accelerate training process and help the emergence of fine details even in the early stage of training. Also, without the term rescaling mechanism, the training process suffers from the scale mismatch between mode-seeking and mode-disengaging terms, which has been pointed out by our analysis, and produces blurred 3D assets. Particularly, a naive combination of the three terms, like SDS, is inefficient and causes blurs.
6 Conclusion
We propose Stable Score Distillation (SSD) for high-quality 3D content generation, based on a comprehensive understanding of the widely used score distillation sampling (SDS). We interpret SDS as a combination of mode-disengaging, mode-seeking and variance-reducing terms, analyzing their distinct properties. This interpretation allows us to harness each term to its fullest potential and to leverage their complementary nature. Our analysis establishes rich connections to prevalent observations and practices in 3D content generation. Extensive experiments demonstrate the effectiveness of our approach for generating high-fidelity 3D content without succumbing to issues such as over-smoothness and severe floaters.
@@ 6 Conclusion
In this supplementary material, Sec. 7 presents more qualitative results of SSD on diverse prompts. We provide proofs in Sec. 8, Sec. 9 and Sec. 10, as well as numerical studies in Sec. 11 to support our method’s assumptions and analysis. Sec. 12 shows the connection between prevalent observations and our analysis. Sec. 13 gives a numerical experiment that shows the efficiency of SSD as a general-purpose mode approximator.
@@ 7 More Qualitative Results
We present additional 3D generation results with more prompts in Fig. 11 and Fig. 12.
Figure 7: Evaluating the learning behaviour of $\theta$ supervised by different estimators with an illustrative 2D example. In this toy example $\theta=\boldsymbol{x}\in\mathbb{R}^{2}$, and $\mathbb{P}(\boldsymbol{x})$ is a mixture of Gaussian distributions. The setup details are in Sec. 13. We initialize $\theta$ to be at the yellow star, optimize it with Eq. 3 under different estimators, and record the learning trajectories of $\theta$. Note that ideally $\theta$ should converge to any modes around the red stars. Mode-Disengaging trajectory: the learning trajectory of $\boldsymbol{x}$ by employing the mode-disengaging term only in Eq. 3, where Mode-Seeking trajectory and SSD trajectory are defined similarly. The purple trajectory describes how $\theta$ evolves when initialized at the trapping point and supervised by mode-disengaging term. On the right, we verify that the mode-seeking term indeed tries to approximate the transient mode for large $t$, causing the convergence point to lie between the normal modes of $\mathbb{P}(\boldsymbol{x})$ on the left. We present the detailed experiment analysis in Sec. 13.
@@
Figure 9: Visualization of variance-reduction schemes for gradients. Note that the SDS-like variance-reduction scheme does not effectively reduce variance of the conditional score for small $t$. When $t$ is large, e.g., 500, the adaptive variance-reduction term, $c$, is close to 1 and our variance reduction scheme performs similarly to SDS-like variance reduction scheme. On the other hand, when $t$ is small, our method significantly differs from SDS and produces smoother gradients. Conditional score represents the gradients got by using the mode-seeking term only in optimization.
@@
Figure 13: Comparison on prompt “a baby dragon drinking boba”. With the proposed adaptive variance reduction scheme, SSD generates clear outline of the boba with only 2000 training steps. The accelerated learning pace on local features facilitates the successful generation of detailed boba, in contrast to the fuzziness observed in the absence of the adaptive variance reduction scheme.
Figure 14: Comparison on prompt “a bald eagle carved out of wood”. With the adaptive variance reduction scheme, SSD learns faster and generates plausible squama and colors of the carved eagle with only 3000 training steps. This scheme effectively alleviates the over-smoothing problem and the final generation shows delicate squama whereas the same region is overly smooth without the scheme.
@@
8 Analysis for Optimal c
For ease of mathematical analysis, in this work we adopt total variance as our variance definition in Eq. 5 for its wide usage [41]. We define $\Sigma_{XY}$ as the covariance matrix between random vectors $X$ and $Y$, and define $\Sigma_{X}$ as a shorthand for $\Sigma_{XX}$. Then Eq. 5 becomes $c(\boldsymbol{x},t)=\text{argmin}_{k}\ \text{tr}(\Sigma_{\hat{\boldsymbol{\epsilon}}_{\phi}-k\boldsymbol{\epsilon}})$, which has a closed-form solution (see Sec. 9 for proof):
Note that $r$ is defined this way to minimize the norm of $\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)-k\boldsymbol{\epsilon}$, namely $r(\boldsymbol{x}_{t},\boldsymbol{\epsilon},y,t)=\text{argmin}_{k}||\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)-k\boldsymbol{\epsilon}||.$ Our experiments reveal that $r$ is quite robust to $\boldsymbol{\epsilon}$ for any combination of $\boldsymbol{x}$, $y$, and $t$, and $c$ is tightly bounded by the extremes of $r$, as shown in App. Fig. 18.
9 Proof for the Closed Form of Optimal c.
@@ 11 More Numerical Experiments
Figure 16: $||\hat{\boldsymbol{\epsilon}}_{\phi}||$ increases as diffusion timestep $t$ increases. This together with Fig. 17 justifies our design of adaptive variance reduction. See Sec. 11 for the data collection protocol.
@@
Figure 17: Linear correlation [25] between $\hat{\boldsymbol{\epsilon}}_{\phi}$ and $\boldsymbol{\epsilon}$ increases with diffusion timestep $t$. This together with Fig. 16 justifies our design of adaptive variance reduction. See Sec. 11 for the data collection protocol.
Figure 18: The optimal $c(\boldsymbol{x},t)$ lies within the extremes of our proposed proxy $r$. For each timestep, we compute $r$ as defined in Eq. 6 and $c$ as defined in Eq. 8. Note that with the 8192 samples as discussed in Sec. 11, $c$ is unique while $r$ can take 8192 different values. Therefore the maximum and minimum of these $r$ are visualized separately. It is evident that $r$ and $c$ are numerically close with similar trend, validating that $r$ is a feasible proxy for $c$. See Sec. 11 for the data collection protocol.
@@
Figure 19: Comparison between norms of the mode-seeking and mode-disengaging terms. Note that the y-axis is plotted in log scale. $||\hat{\boldsymbol{\epsilon}}_{\phi}||\in[80,130]$ while $||h||\in[2,5.5]$. The vast norm difference between the two terms constitutes an important justification for our estimator-rescaling in the proposed stable score distillation (SSD) Algorithm 1 (L15). See Sec. 11 for the data collection protocol.
12 Connection with Common Observations and Practices in 3D Content Generation
  • Large CFG Scales. As illustrated in App. Fig. 19, the norms of the mode-disengaging term and the mode-seeking term are vastly different. Consequently, a large CFG scale is necessary to make the scale of the main learning signal $\omega\cdot h$ at least comparable to that of the mode-seeking term. Otherwise the SDS estimator in Eq. 3 would be dominated by the mode-seeking term, which can only generate over-smoothed contents for reasons discussed in Sec. 4.3.2, as visualized in Fig. 10.
  • Color Saturation. This challenge has been repeatedly observed in text-to-3D generation [24]. Recent work of DreamTime [11] also observes severe color saturation tendency when $t<100$. According to our analysis in Sec. 4.2, when $t$ is small the mode-disengaging term tends to drive the rendering $g(\theta,c)$ away from modes of the natural image distribution, making the rendering implausible. As pointed out in [4], in practice the diffusion model tends to influence fine details and colors, but not large-scale geometry, when $t$ is small. Thus with small $t$ the mode-disengaging term generates implausible colors, without corrupting the general geometry of the objects. We provide experiments in Fig. 20 to show that the mode-disengaging term actually causes color saturation with 2D experiments.
Figure 20: Comparisons on 2D optimization. The experiments are run with more training iterations than normally used to expose the intrinsic properties of the estimators. Objects’ local features are more plausible with SSD. It is evident that while the mode-seeking term contributes to finer local details, it also incurs color implausibility.
13 Illustrative Example
We evaluate the learning behaviour of the mode-seeking, mode-disengaging and SSD terms with an illustrative example, as shown in Fig. 7. In this experiment, we assume that $\boldsymbol{x}=\theta\in\mathbb{R}^{2}$, and $\mathbb{P}_{\phi}(\boldsymbol{x})=0.2\mathcal{N}([0,0]^{T},0.1I)+0.4\mathcal{N}([1,1]^{T},0.05I)+0.4\mathcal{N}([2,1]^{T},0.05I)$ is a mixture of Gaussian distributions. The two Gaussian components located at $[1,1]^{T}$ and $[2,1]^{T}$ are conditional modes, while the one at $[0,0]^{T}$ is a singular one. Note that we want $\theta$ to approximate the two conditional modes as close as possible. We initialize $\theta$ at $[0,0]^{T}$ to simulate the situation that all 3D generation algorithms begin with empty renderings. We then substitute the mode-seeking, mode-disengaging and our SSD estimators for the SDS estimator in Eq. 3 one by one, and record the learning trajectory of $\theta$. To validate the trap-escaping property of the mode-disengaging term we initialize a $\theta$ to the trapping point, namely $[1.5,1]^{T}$, and use $h$ to supervise $\theta$ afterwards. The learning trajectories are illustrated in Fig. 7 (left). We observe that the mode-seeking trajectory gets trapped between the two conditional modes rapidly, not converging to any specific conditional mode of $\mathbb{P}_{\phi}(\boldsymbol{x})$. To inspect whether the issue is caused by the transient-mode problem, we present the density map of $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y)$ at $t=350$, and the learning trajectory of $\alpha_{t}\boldsymbol{x}$ in Fig. 7 (right). The density map reveals that the two conditional modes of $\mathbb{P}_{\phi}(\boldsymbol{x};y)$ result in a transient mode in $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y)$, whose probability density is higher than the induced modes. And $\alpha_{t}\boldsymbol{x}_{t}$ indeed approaches the transient mode $\alpha_{t}\boldsymbol{o}_{tr}$. Conversely, the mode-disengaging term can propel $\theta$ away from the trapping point $\boldsymbol{o}_{tr}$. However, although at the beginning it guides $\theta$ towards a conditional mode, it ultimately steers $\theta$ into a low-density region of $\mathbb{P}_{\phi}(\boldsymbol{x})$. In contrast, the proposed SSD swiftly guides $\theta$ towards a conditional mode at beginning, avoids getting trapped, and finally converges to a point with high density in $\mathbb{P}_{\phi}(\boldsymbol{x})$.